Rai Corporation had 250,000 shares of common stock outstanding during the current year. There were also fully vested options for 10,500 shares of common stock were granted with an exercise price of $20. The market price of the common stock averaged $25 for the year. Net income was $4.5 million. What is diluted EPS (rounded)?
Answer #1

10,500-(10,500*20/25)= 2,100

250,000+2,100= 252,100

4,500,000/252,100= 17.8500595002 or 17.85 answer

answered by: anonymous